let 'f' be a continuous function on 'R' such that


f(1/2n)=(sin(en))e-n^2+2[n]/n2+[n]+1 where [.] denotes greatest integer function. Then find the value of f(0).

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Paidepelly

f(1/2n)=(sin(en))e-n^2+2[n]/(n2+[n]+1)

for f(0) take limit n tends to infinity both side

Lt n→∞ f(1/2n)=Lt n→∞(sin(en))e-n^2+2[n]/(n2+[n]+1)

f(0)=Lt n→∞(sin(en))e-n^2+2[n]/(n2+[n]+1)

        =Lt n→∞(sin(en))e-n^2+ Lt n→∞   2[n]/(n2+[n]+1)

        =0  +Lt n→∞   2[n]/(n2+[n]+1)

let n=[n] +f wher f is fractional part of n

so f(0) =Lt n→∞   2(n-f)/(n2+n-f+1)

find limit

f(0) =2


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin